### v1

@@ 4 Policy iteration theorem & RL algorithm
Throughout this section, we fix a sufficiently large $N\in\mathbb{N}$ and a small $\lambda>0$ so that $\overline{Y}^{x;N,\lambda}$ serves as an accurate approximation of $Y^{x}$ (see Remark 2.9 and Theorem 3.5). The proofs of all theorems in this section can be found in Section 6.3.
@@
(4.1) $\pi^{n+1}_{t}:=\operatorname{logit}(\frac{N}{\lambda}(R(X_{t}^{x})-\overline{Y}_{t}^{x;N,\lambda,\pi^{n}})),\quad t\in[0,T].$
@@
    $\overline{Y}_{t}^{x;N,\lambda}\geq\overline{Y}_{t}^{x;N,\lambda,\pi^{n+1}}\geq\overline{Y}_{t}^{x;N,\lambda,\pi^{n}}$, $\mathbb{P}$-a.s., for all $t\in[0,T]$;
@@
In particular, $\overline{Y}^{x;N,\lambda,\pi^{n}}_{t}\uparrow\overline{Y}^{x;N,\lambda}_{t}$ and $\pi^{n}_{t}\uparrow\pi^{*}_{t}$ $\mathbb{P}$-a.s. for all $t\in[0,T]$ as $n\to\infty$.
@@
Clearly, $\check{F}_{t}^{N,\lambda}(X_{t}^{x},y,z)=\overline{F}^{x;N,\lambda}_{t}(y,z)$ for $(t,x,y,z)\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{d}$; see (3.7). Moreover, $\check{F}_{\cdot}^{N,\lambda}(\cdot,\cdot,\cdot)$ and $R(\cdot)$ satisfy the conditions (M1b) and ($\textrm{M1b}^{c}$) given in [19]. Therefore, an application of [19, Theorem 8.12] ensures the existence of a viscosity solution [^2] We refer to [19, Definition 8.11] for the definition of a viscosity solution of (4.3) with setting the terminal condition $R\curvearrowright\Psi$ and the generator $\check{F}^{N,\lambda}_{\cdot}\curvearrowright g$ therein. $\check{v}^{N,\lambda}$ of the following PDE:,
(4.3) $(\partial_{t}v+\mathcal{L}v)(t,x)+\check{F}^{N,\lambda}_{t}\big(x,v(t,x),((\widetilde{\sigma}^{o})^{\top}\nabla v)(t,x)\big)=0\;\;\;(t,x)\in[0,T)\times\mathbb{R}^{d},$
@@
We now have a sequence of closed-loop policies in $\check{\Pi}$ deriving the policy iteration.
@@
    $\overline{Y}_{t}^{x;N,\lambda,\check{\pi}^{n}(X^{x})}=v^{N,\lambda,n}(t,X_{t}^{x}),\qquad\overline{Z}_{s}^{x;N,\lambda,\check{\pi}^{n}(X^{x})}=\big((\widetilde{\sigma}^{o})^{\top}w^{N,\lambda,n}\big)(t,X_{t}^{x}),$
@@
    $(\partial_{t}v+\mathcal{L}v)(t,x)+\check{F}^{N,\lambda,\check{\pi}^{n}}_{t}(x,v(t,x),((\widetilde{\sigma}^{o})^{\top}\nabla v)(t,x))=0\;\;\;(t,x)\in[0,T)\times\mathbb{R}^{d},$
@@
In what follows, we develop an RL scheme, relying on the deep splitting method of Beck et al. [5] and Frey and Köck [25], to implement the policy evaluation step at each iteration. For this purpose, we first introduce some notation, omitting the dependence on $(N,\lambda)$ (even though the objects still depend on them).
@@
    $\begin{aligned} \check{F}_{i}^{n}(\tilde{x},y,z;\vartheta_{i}^{n-1})&:=r(\tilde{x})-\beta_{t_{i}}y+g(t,z)+N(R(\tilde{x})-y)\check{\pi}_{i}^{n}(\tilde{x};\vartheta_{i}^{n-1})\\ &\qquad-\lambda\mathcal{H}\big(\check{\pi}_{i}^{n}(\tilde{x};\vartheta_{i}^{n-1})\big),\end{aligned}$
@@
so that $\Sigma({\check{X}^{x}_{i:i+1}})\Sigma({\check{X}^{x}_{i:i+1}})^{\top}\Delta t\to\widetilde{\sigma}^{o}(t_{i},\check{X}^{x}_{i})\widetilde{\sigma}^{o}(t_{i},\check{X}^{x}_{i})^{\top}\Delta t$ as $\Delta t\downarrow 0$ in probability $\mathbb{P}$; see e.g., [34, Chapter I, Theorem 4.47] and [56, Section 6, Theorem 22].
@@
With all this notation set in place, for each iteration $n\in\mathbb{N}$, we present the policy evaluation as the following iterative minimization problem: for $i\in\{0,\dots,I-1\}$
@@
where $\mathfrak{L}_{i}^{n}(\cdot;\vartheta_{i}^{n-1,*},\vartheta_{i+1}^{n,*}):\Theta\to\mathbb{R}$ is the (parameterized) $L^{2}$-loss function given by
@@
(4.6) $\quad\quad+\check{F}^{n}_{i}\big(\check{X}^{x}_{i+1},v^{n}_{i+1}(\check{X}^{x}_{i+1};\theta^{n,*}_{i+1}),\Sigma({\check{X}^{x}_{i:i+1}})\nabla v^{n}_{i+1}(\check{X}^{x}_{i+1};\theta^{n,*}_{i+1});\vartheta_{i}^{n-1,*}\big)\Delta t\big|^{2}\Big],$
@@
We numerically solve the problem given in (4.5) by using stochastic gradient descent (SGD) algorithms (see, e.g., [28, Section 4.3]). Then we provide a pseudo-code in Algorithm 1 to show how the policy iteration can be implemented.
@@
Note that the deep splitting method of [5, 25] is not the only neural realization of our policy evaluation; instead deep BSDEs / PDEs schemes of [30, 33, 62] can be an alternative. More recently, several articles, including [27, 46], provide the error analyses for such methods. To obtain a full error-analysis of our policy iteration algorithm, one would need to relax the standard Lipschitz and Hölder conditions on BSDE generators in the mentioned articles so as to cover the generator $\check{F}^{N,\lambda,\check{\pi}^{n}}$ in (4.2), and then incorporate the policy evaluation errors from the neural approximations (under such relaxed conditions) into the convergence rate established in Theorem 4.1. We defer this direction to a future work.
@@ 5 Experiments
We conduct experiments on the American put and call holder’s stopping problems to illustrate the policy improvement, convergence, stability, and robustness of Algorithm 1. The simulation settings are as follows: under Setting 4, we let the running reward $r(\cdot)\equiv 0$, the discounting factor $\beta_{t}\equiv r_{*}$, the volatility $\widetilde{\sigma}^{o}(t,\check{x})=0.4\check{x}$, the initial price and strike price $x=\Gamma=40$, and
@@
We first examine the policy improvement and convergence of Algorithm 1. For the put-type stopping problem, we fix $\lambda=1$ and $N=10$, and consider several ambiguity degrees $\varepsilon\in\{0,0.2,0.4\}$. The reference values $R^{\mathrm{ref}}_{\varepsilon}$ for $\varepsilon\in\{0,0.2,0.4\}$ are obtained by solving the BSDE (3.8) for the corresponding optimal value function using the deep backward scheme of Huré et al. [33], yielding $R^{\mathrm{ref}}_{0}=5.302$, $R^{\mathrm{ref}}_{0.2}=4.420$, $R^{\mathrm{ref}}_{0.4}=3.725$. The results illustrating the policy improvement and convergence are shown in Figure 1, which align well with the theoretical findings in Theorem 4.1.
@@
Figure 1: Policy improvement and convergence in Algorithm 1 under several ambiguity levels.
To examine the stability of Algorithm 1, we vary the penalty, temperature and ambiguity levels as $N\in\{5,10,20\}$, $\lambda\in\{0.01,1,5\}$, and $\varepsilon\in\{0,0.2,0.4\}$, and present the corresponding values of $v^{N,\lambda,\star;\varepsilon}_{0}$ in Table 1 (obtained after at-least 10 iterations of the policy improvement; see Figure 1). These results align with the stability analysis w.r.t. $\lambda$ given in Theorem 3.5 and the sensitivity analysis of robust optimization problems w.r.t. ambiguity level examined in [2, Theorem 2.13], [10, Corollary 5.4].
Table 1: Stability analysis of Algorithm 1 w.r.t. the penalty, temperature and ambiguity levels.
@@
Figure 2: Robustness performance under unknown testing environments.
@@
Lastly, we examine the robustness of Algorithm 1 in the call-type stopping problem. In particular, to assess the out-of-sample performance under an unknown testing environment, we re-simulate new state trajectories $(\check{X}^{x,\delta}_{i})_{i=0}^{I}$ as in Setting 4 (i) under different dividend rates $\delta\in\{0,0.05,0.1,0.15,0.2,0.25\}$, where the number of simulated trajectories is set to $2^{20}$. We fix $N=10$ and consider configuration $\varepsilon\in\{0,0.1,0.2,0.3\}$ both for $\lambda=1$ and $\lambda=5$. Using the trained value functions $(v^{10,\lambda,\star;\varepsilon}_{i}(\cdot))_{i=0}^{I}$, the stopping policy $\tau_{\delta}^{\varepsilon,\lambda}$ and corresponding discounted expected reward $\check{R}^{\varepsilon,\lambda}_{\delta}$ under such unknown environment are defined by
$\tau^{\varepsilon,\lambda}_{\delta}$ $:=\inf\big\{t_{i}:v^{10,\lambda,\star;\varepsilon}_{i}(\check{X}^{x,\delta}_{i})\leq R(\check{X}^{x,\delta}_{i}),\;i=0,\ldots,I\big\},$
$\check{R}^{\varepsilon,\lambda}_{\delta}$ $:=\mathbb{E}\big[e^{-r_{*}\tau^{\varepsilon,\lambda}_{\delta}}R(\check{X}^{x,\delta}_{i})\big].$
@@
For each $\delta$, the corresponding American call option price represents the optimal value for the call-type stopping problem, which can be computed using the implicit finite-difference method of Forsyth and Vetzal [24]. We therefore use the option prices computed by this method as reference values $R^{\mathrm{ref}}_{\delta}$ for each $\delta$, yielding $R^{\mathrm{ref}}_{0}=4.954,$ $R^{\mathrm{ref}}_{0.05}=4.410$, $R^{\mathrm{ref}}_{0.1}=3.990$, $R^{\mathrm{ref}}_{0.15}=3.634$, $R^{\mathrm{ref}}_{2}=3.324$, $R^{\mathrm{ref}}_{0.25}=3.052$. The relative errors are then computed as ${|\check{R}^{\varepsilon,\lambda}_{\delta}-R^{\mathrm{ref}}_{\delta}|}/{R^{\mathrm{ref}}_{\delta}}$.
@@
In Figure 2, when the dividend rate in the testing environment does not deviate significantly from that of the trained environment (near $\delta=0.05$), the non-robust value function (i.e., with $\varepsilon=0$) performs comparably well. However, as the discrepancy between the training and testing environments increases, the benefit of incorporating ambiguity into the framework becomes evident, as reflected by lower relative errors for higher ambiguity levels (e.g., $\varepsilon=0.2,0.3$).

### v2

@@ 4 Policy iteration theorem & RL algorithm
Throughout this section, we fix a sufficiently large $N\in\mathbb{N}$ and a small $\lambda>0$ so that $\overline{Y}^{x;N,\lambda}$ serves as an approximation of $Y^{x}$ (see Remark 2.9 and Theorem 3.5).
@@
(4.1) $\pi^{n+1}_{t}:=\operatorname{logit}\bigg(\frac{N}{\lambda}(R(X_{t}^{x})-\overline{Y}_{t}^{x;N,\lambda,\pi^{n}})\bigg),\quad t\in[0,T].$
@@
    $\overline{Y}_{t}^{x;N,\lambda}\geq\overline{Y}_{t}^{x;N,\lambda,\pi^{n+1}}\geq\overline{Y}_{t}^{x;N,\lambda,\pi^{n}}$, $t\in[0,T]$, $\mathbb{P}$-a.s.;
@@
In particular, for every $t\in[0,T]$, $\overline{Y}^{x;N,\lambda,\pi^{n}}_{t}\uparrow\overline{Y}^{x;N,\lambda}_{t}$ and $\pi^{n}_{t}\downarrow\pi^{*}_{t}$ $\mathbb{P}$-a.s., as $n\to\infty$.
The proof of Theorem 4.1 can be found in Section 6.3.
Remark 4.2.
The policy iteration result in [dai2024learning, Theorem 4] builds on PDE arguments in a Markovian framework, whereas Theorem 4.1 is based on BSDE arguments and shows convergence rates in a non-Markovian setting. In particular, [dai2024learning] considers a bounded payoff and an initial iterate value function that is continuous and bounded, whereas our result applies to payoffs without boundedness condition and does not require an initial value function; instead, the iteration is induced by any initial policy.
A BSDE-based policy iteration result for Dynkin games is studied in [DaiDong2024-dynkin, Theorem 3] under boundedness conditions on their BSDE paramters. While the iterative arguments using the comparison principle of BSDEs therein might be similar to ours, the presence of the generator $\overline{Z}^{x;N,\lambda,\pi^{n}}$ in our BSDE driver, induced by ambiguity, requires a more delicate analysis of a priori estimates and iterative arguments. In particular, our approach does not require boundedness of our BSDE parameters.
Remark 4.3.
Let $N\in\mathbb{N}$ and $\lambda>0$ be given. Using $(\pi^{n})_{n\in\mathbb{N}}$ in (4.1), consider a sequence of stopping times $(\tau^{x;N,\lambda,n}_{t})_{n\in\mathbb{N}}\subset{\mathcal{}T}_{t}$ defined, for every $n\in\mathbb{N}$, by
$\tau^{x;N,\lambda,n}_{t}:=$ $\inf\bigg\{s\geq t:\pi^{n+1}_{s}\geq\frac{1}{1+2^{N(T-s)}}\bigg\}\wedge T$
$=$ $\inf\{s\geq t:\overline{Y}_{s}^{x;N,\lambda,\pi^{n}}-\lambda(T-s)\log 2\leq R(X_{s}^{x})\}\wedge T\quad t\in[0,T].$
From Theorem 4.1 and definition of $\tau^{x;N,\lambda}_{t}$ in (3.12), we can use the same arguments presented for Theorem 3.6 to have that for any $t\in[0,T]$, $\tau^{x;N,\lambda,n}_{t}\to\tau^{x;N,\lambda}_{t}$ $\mathbb{P}$-a.s., as $n\to\infty$. Moreover, for any $t\in[0,T]$, $\|{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{x;N,\lambda,n}}]-{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{x;N,\lambda}}]\|_{L^{2}}\to 0$ as $n\to\infty$.
@@
Clearly, $\check{F}_{t}^{N,\lambda}(X_{t}^{x},y,z)=\overline{F}^{x;N,\lambda}_{t}(y,z)$ for $(t,x,y,z)\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{R}^{d}$; see (3.7). Moreover, $\check{F}_{\cdot}^{N,\lambda}(\cdot,\cdot,\cdot)$ and $R(\cdot)$ satisfy the conditions (M1b) and ($\textrm{M1b}^{c}$) given in [KHM]. Therefore, an application of [KHM, Theorem 8.12] ensures the existence of a viscosity solution [^1] We refer to [KHM, Definition 8.11] for the definition of a viscosity solution of (4.3) with setting the terminal condition $R\curvearrowright\Psi$ and the generator $\check{F}^{N,\lambda}_{\cdot}\curvearrowright g$ therein. $\check{v}^{N,\lambda}$ of the following PDE:,
(4.3) $(\partial_{t}v+\mathcal{L}v)(t,x)+\check{F}^{N,\lambda}_{t}\big(x,v(t,x),((\widetilde{\sigma}^{o})^{\top}\nabla v)(t,x)\big)=0,\;\;\;(t,x)\in[0,T)\times\mathbb{R}^{d},$
@@
We now have a sequence of closed-loop policies in $\check{\Pi}$ deriving the policy iteration. The proof is presented in Section 6.3.
@@
    $\overline{Y}_{t}^{x;N,\lambda,\check{\pi}^{n}(X^{x})}=v^{N,\lambda,n}(t,X_{t}^{x}),\quad\overline{Z}_{s}^{x;N,\lambda,\check{\pi}^{n}(X^{x})}=\big((\widetilde{\sigma}^{o})^{\top}w^{N,\lambda,n}\big)(t,X_{t}^{x}),$
@@
    $(\partial_{t}v+\mathcal{L}v)(t,x)+\check{F}^{N,\lambda,\check{\pi}^{n}}_{t}(x,v(t,x),((\widetilde{\sigma}^{o})^{\top}\nabla v)(t,x))=0,\;\;(t,x)\in[0,T)\times\mathbb{R}^{d},$
@@
Remark 4.6.
Corollary 4.5 is a closed-loop, Markovian counterpart to the open-loop policy iteration in Theorem 4.1. In particular, measurability and continuity on the initial policy ensure that the induced sequence of policies admits feedback-form representations in (4.4). As noted in Remark 4.2, [dai2024learning, Theorem 4] establishes a PDE-based policy iteration result. However, our result does not follow from it, since we do not assume boundedness of the reward functions and the associated iterative PDEs lack Lipschitz regularity in source terms. Moreover, the presence of $\overline{Z}^{x;N,\lambda,\check{\pi}^{n}(X^{x})}$, induced by ambiguity, requires additional analysis. Our result is therefore obtained via a direct probabilistic approach, based on [KHM, Theorem 8.12].
@@
In what follows, we develop an RL scheme, relying on the deep splitting method of Beck et al. [Becker-Jentzen-Neufeld-Deep-Splitting-21] and Frey and Köck [deep_splitting_convergence], to implement the policy evaluation step at each iteration. For this purpose, we first introduce some notation, omitting the dependence on $(N,\lambda)$ (even though the objects still depend on them).
@@
    $\begin{aligned} \check{F}_{i}^{n}(\tilde{x},y,z;\vartheta_{i}^{n-1})&:=r(\tilde{x})-\beta_{t_{i}}y+g(t_{i},z)+N(R(\tilde{x})-y)\check{\pi}_{i}^{n}(\tilde{x};\vartheta_{i}^{n-1})\\ &\quad-\lambda\mathcal{H}\big(\check{\pi}_{i}^{n}(\tilde{x};\vartheta_{i}^{n-1})\big),\end{aligned}$
@@
so that $\Sigma({\check{X}^{x}_{i:i+1}})\Sigma({\check{X}^{x}_{i:i+1}})^{\top}\Delta t\to\widetilde{\sigma}^{o}(t_{i},\check{X}^{x}_{i})\widetilde{\sigma}^{o}(t_{i},\check{X}^{x}_{i})^{\top}\Delta t$ as $\Delta t\downarrow 0$ in probability $\mathbb{P}$; see e.g., [jacod2013limit, Chapter I, Theorem 4.47] and [Protter2005Stochastic, Section 6, Theorem 22].
@@
With all this notation set in place, for each iteration $n\in\mathbb{N}$, we present the policy evaluation as a backward iterative minimization problem. For any $i\in\{0,\dots,I-1\}$, let $\vartheta^{n-1,*}_{i},\;i\in\{0,\ldots,I-1\}$ denote the optimal parameter at time $t_{i}$ from iteration n1n-1, and let $\vartheta^{n,*}_{i+1}$ denote the optimal parameter at time $t_{i+1}$ in $n$-iteration. Then the optimal parameter at time $t_{i}$ in $n$-iteration is given by
@@
where $\mathfrak{L}_{i}^{n}(\cdot;\vartheta_{i}^{n-1,*},\vartheta_{i+1}^{n,*}):\Theta\to\mathbb{R}$ is the parameterized $L^{2}$-loss function given by
@@
(4.6) $\quad\quad+\check{F}^{n}_{i}\big(\check{X}^{x}_{i+1},v^{n}_{i+1}(\check{X}^{x}_{i+1};\vartheta^{n,*}_{i+1}),\Sigma({\check{X}^{x}_{i:i+1}})\nabla v^{n}_{i+1}(\check{X}^{x}_{i+1};\vartheta^{n,*}_{i+1});\vartheta_{i}^{n-1,*}\big)\Delta t\big|^{2}\Big],$
@@
We numerically solve the problem given in (4.5) by using stochastic gradient descent (SGD) algorithms (see, e.g., [goodfellow2016deep, Section 4.3]). Then we provide a pseudo-code in Algorithm 1 to show how the policy iteration can be implemented.
@@
Note that the deep splitting method of [Becker-Jentzen-Neufeld-Deep-Splitting-21, deep_splitting_convergence] is not the only neural realization of our policy evaluation; instead deep BSDEs / PDEs schemes of [han18, pham2020deepBSDE, sirignano2018dgm] can be an alternative. More recently, several articles, including [pham2022deepBSDE_erroranalysis, neufeld2024full], provide the error analyses for such methods. To obtain a full error-analysis of our policy iteration algorithm, one would need to relax the standard Lipschitz and Hölder conditions on BSDE generators in the mentioned articles so as to cover the generator $\check{F}^{N,\lambda,\check{\pi}^{n}}$ in (4.2), and then incorporate the policy evaluation errors from the neural approximations (under such relaxed conditions) into the convergence rate established in Theorem 4.1. We defer this direction to a future work.
@@ 5 Experiments
5.1 Example 1
We first conduct experiments on 1-dimensional American put / call holder’s stopping problems to illustrate the policy improvement, convergence, stability, and robustness of Algorithm 1. The simulation settings are as follows: under Setting 4, we let the running reward $r(\cdot)\equiv 0$, the discounting factor $\beta_{t}\equiv r_{*}$, the volatility $\widetilde{\sigma}^{o}(t,x)=0.4x$, $x\in\mathbb{R}$, the initial price and strike price $x_{0}=\Gamma=40$, and
@@ 5.1 Example 1
We first examine the policy improvement and convergence of Algorithm 1. For the put-type stopping problem, we fix $\lambda=1$ and $N=10$, and consider several ambiguity degrees $\varepsilon\in\{0,0.2,0.4\}$. The reference values $R^{\mathrm{ref}}_{\varepsilon}$ for $\varepsilon\in\{0,0.2,0.4\}$ are obtained by solving (3.8) for the corresponding optimal value function using the deep backward scheme of Huré et al. [pham2020deepBSDE], yielding $R^{\mathrm{ref}}_{0}=5.302$, $R^{\mathrm{ref}}_{0.2}=4.420$, $R^{\mathrm{ref}}_{0.4}=3.725$. The results illustrating the policy improvement and convergence are shown in Fig. 1, which align well with the theoretical findings in Theorem 4.1.
@@
Figure 1: Policy improvement and convergence of Algorithm 1 in Example 1.
To examine the stability of Algorithm 1, we vary the penalty, temperature and ambiguity levels as $N\in\{5,10,20\}$, $\lambda\in\{0.01,1,5\}$, and $\varepsilon\in\{0,0.2,0.4\}$, and present the corresponding values of $v^{N,\lambda,\star;\varepsilon}_{0}$ in Table 1 (obtained after at-least 10 iterations of the policy improvement; see Fig. 1). These results align with the stability analysis w.r.t. $\lambda$ given in Theorem 3.5 and the sensitivity analysis of robust optimization problems w.r.t. ambiguity level examined in [bartl2023sensitivity, Theorem 2.13], [chen2024robust, Corollary 5.4].
Table 1: Stability of Algorithm 1 w.r.t. the penalty, temperature and ambiguity levels in Example 1.
@@
Figure 2: Robustness performance under unknown testing environments in Example 1.
@@
Lastly, we examine the robustness of Algorithm 1 in the call-type stopping problem. In particular, to assess the out-of-sample performance under an unknown testing environment, we re-simulate new state trajectories $(\check{X}^{x,\delta}_{i})_{i=0}^{I}$ as in Setting 4 (i) under different dividend rates $\delta\in\{0,0.05,0.1,0.15,0.2,0.25\}$, where the number of simulated trajectories is set to $2^{20}$. We fix $N=10$ and consider configuration $\varepsilon\in\{0,0.1,0.2,0.3\}$ both for $\lambda\in\{0.01,1,5\}$. Using the trained value functions $(v^{10,\lambda,\star;\varepsilon}_{i}(\cdot))_{i=0}^{I}$, the stopping time $\tau_{\delta}^{\varepsilon,\lambda}$ and the corresponding discounted expected reward $\check{R}^{\varepsilon,\lambda}_{\delta}$ under such unknown environment are defined by
$\tau^{\varepsilon,\lambda}_{\delta}$ $:=\inf\big\{t_{i}:v^{10,\lambda,\star;\varepsilon}_{i}(\check{X}^{x,\delta}_{i})-{\bf 1}_{\{\lambda=0.01\}}\lambda(T-t_{i})\log 2\leq R(\check{X}^{x,\delta}_{i}),\;i=0,\ldots,I\big\},$
(5.1) $\quad\check{R}^{\varepsilon,\lambda}_{\delta}:=\mathbb{E}[e^{-r_{*}\tau^{\varepsilon,\lambda}_{\delta}}R(\check{X}^{x,\delta}_{i})].$
@@
For each $\delta$, the corresponding American call option price represents the optimal value for the call-type stopping problem, which can be computed using the implicit finite-difference method of Forsyth and Vetzal [Forsyth-Penalty-02]. We therefore use the option prices computed by this method as reference values $R^{\mathrm{ref}}_{\delta}$ for each $\delta$, yielding $R^{\mathrm{ref}}_{0}=4.954,$ $R^{\mathrm{ref}}_{0.05}=4.410$, $R^{\mathrm{ref}}_{0.1}=3.990$, $R^{\mathrm{ref}}_{0.15}=3.634$, $R^{\mathrm{ref}}_{2}=3.324$, $R^{\mathrm{ref}}_{0.25}=3.052$. The relative errors are then computed as ${|\check{R}^{\varepsilon,\lambda}_{\delta}-R^{\mathrm{ref}}_{\delta}|}/{R^{\mathrm{ref}}_{\delta}}$.
Moreover, we note that the stopping rule in (5.1) follows the robust exploratory stopping time $\tau_{0}^{x;N,\lambda}$ in (3.12) for the case $\lambda=0.01$, because it approximates the optimal stopping time $\tau^{*,x}_{0}$ in (2.5) for sufficiently small $\lambda$ and large $N$ (see Theorem 3.6). For larger values of $\lambda$, in particular $\lambda=1,5$, we heuristically adopt the stopping rule $\overline{\tau}_{0}^{x;N,\lambda}$ in Remark 3.7 without the adjustment term $\lambda(T-t_{i})\log 2$, since the adjustment term becomes relatively large and may not only impact the approximation accuracy, but also diminish the exploration benefit.
@@
Last, we note in Fig. 2 that when the dividend rate in the testing environment does not deviate significantly from that of the trained environment (near $\delta=0.05$), the non-robust value function (i.e., with $\varepsilon=0$) performs comparably well. However, as the discrepancy between the training and testing environments increases, the benefit of the robust framework becomes evident, as reflected by lower relative errors for higher ambiguity levels (i.e., $\varepsilon=0.2,0.3$). Notably, as the exploration parameter $\lambda$ increases, the robust framework performs better under larger discrepancies between the training and test environments.
5.2 Example 2
Figure 3: Policy improvement and convergence of Algorithm 1 in Example 2
We now implement Algorithm 1 in multi-dimensional settings with $d\in\{5,10\}$. To examine policy improvement and convergence, we consider the American basket put stopping problem. The simulation settings are as follow: $T=1,I=50,\lambda=1$, and $N=10$. The discount factor is $\beta_{t}\equiv 0.06$. The simulator is specified by the volatility and drift functions $\widetilde{\sigma}^{o}(t,x)=\operatorname{diag}(0.4)x$ and $\widetilde{b}^{o}(t,x)=0.06x$, respectively, for $x\in\mathbb{R}^{d}$, where $\operatorname{diag}(a)$ denotes the diagonal matrix with entries $a\in\mathbb{R}$. We consider ambiguity levels $\varepsilon\in\{0,0.2,0.4\}$. Moreover, we set the running reward to $r(\cdot)\equiv 0$, the initial price to $x_{0}=(40,\dots,40)^{\top}\in\mathbb{R}^{d}$, and the strike price to $\Gamma=40$. Last, the payoff is
$R(x)=\mbox{$(\Gamma-\frac{1}{d}\sum_{i=1}^{d}x_{i})^{+},\quad x=(x_{1},\ldots,x_{d})^{\top}\in\mathbb{R}^{d}$}.$
As in Example 1, we compute the reference values $R^{\mathrm{ref}}_{\varepsilon}$ for $\varepsilon\in\{0,0.2,0.4\}$ using the deep backward scheme of [pham2020deepBSDE]. For $d=5$, we obtain $R^{\mathrm{ref}}_{0}=2.086$, $R^{\mathrm{ref}}_{0.2}=1.750$, and $R^{\mathrm{ref}}_{0.4}=1.468$. For $d=10$, we obtain $R^{\mathrm{ref}}_{0}=1.337$, $R^{\mathrm{ref}}_{0.2}=1.105$, and $R^{\mathrm{ref}}_{0.4}=0.947$. The results in Fig. 3 illustrate policy improvement and convergence in high-dimensional settings, supporting the scalability of our algorithm.
Figure 4: Robustness Performance under unknown testing environments in Example 2.
Next, for the robustness test, we consider the geometric call stopping problem with $d=10$. The simulation settings are as follow: $T=0.5$, $I=100$, $\lambda\in\{0.01,1\}$, and $N=10$. The interest rate is specified by $r_{*}=0.05$ and the ambiguity level varies over $\varepsilon\in\{0,0.2,0.4\}$. The drift and volatility functions are specified by $\widetilde{b}^{o}(t,x)=(r_{*}-\delta)x$ and $\widetilde{\sigma}^{o}(t,x)=\operatorname{diag}(0.2)x$, $x\in\mathbb{R}^{d}$, respectively, where we let the dividend rate $\delta$ of the training simulator be $\delta_{\mathrm{train}}=0.05$, whereas we let the dividend rate in the testing simulator vary over $\delta\in\{0,0.05,0.1,0.15,0.2,0.25\}$. Last, the payoff is given by
$R(x)=\big((\prod_{i=1}^{d}x_{i})^{1/d}-\Gamma\big)^{+},\quad x=(x_{1},\ldots,x_{d})^{\top}\in\mathbb{R}^{d}.$
For the evaluation test, we adopt the stopping rule $\tau^{\varepsilon,\lambda}_{\delta}$ and the expected reward $\check{R}_{\delta}^{\varepsilon,\lambda}$ from (5.1), with the payoff replaced by the geometric call function above.
Here we note that, due to the log-normal structure and the assumption of identical drift and volatility across dimensions, the 10-dimensional geometric call problem is equivalent to the one-dimensional problem with $\widetilde{b}^{o}(t,x)=(r_{*}-\delta-\frac{(0.2)^{2}}{2}(1-\frac{1}{d}))x$ and $\widetilde{\sigma}^{o}(t,x)=\frac{0.2}{\sqrt{d}}x$, $x\in\mathbb{R}.$ Hence, we can apply the implicit finite-difference method of Forsyth and Vetzal [Forsyth-Penalty-02] to compute the reference values $R^{\mathrm{ref}}_{\delta}$ in the one-dimensional setting for each $\delta$. The corresponding reference values are given by $R^{\mathrm{ref}}_{0}=1.061$, $R^{\mathrm{ref}}_{0.05}=0.570$, $R^{\mathrm{ref}}_{0.1}=0.350$, $R^{\mathrm{ref}}_{0.15}=0.236$, $R^{\mathrm{ref}}_{0.2}=0.172$, and $R^{\mathrm{ref}}_{0.25}=0.134$. The relative errors are defined by $|\check{R}_{\delta}^{\varepsilon,\lambda}-R^{\mathrm{ref}}_{\delta}|/R^{\mathrm{ref}}_{\delta}$. Fig. 4 illustrates the robustness performance. As in Example 1, our robust RL algorithm performs better as the discrepancy between the training and testing environments increases.